OpenStudy (allyson):

how do you do 5x+7y=3 in form of y+mx+b?

OpenStudy (saifoo.khan):

y=mx + b form or y+mx +b??

OpenStudy (allyson):

sorry y=mx+b lol

OpenStudy (anonymous):

7y=-5x+3 y=(-5x+3)/7 y=-5x/7 +3/7

OpenStudy (anonymous):

\[y=-\frac{5}{7}x+\frac{3}{7}\]

OpenStudy (allyson):

thank you soooo muchh (:

OpenStudy (anonymous):

ur welcome